Chair: Jasper

Next meeting chair: D Smith (28/04)

Attended: Simin, Lim, Ahmad, Earl, Eloise, Jay, Jamie, D Schunck.

Apologies:

General Issues:

  • Ke status, main problems fixed but phase cal still not operational. Now good for mixed mode and auv sessions. Turned out to be and Issue with the dehydrator.
  • Reminder of DST change

Notes:

Another Vgos test coming up next Thursday.

Handover issues:

Yg:

  • R41046: Had issue with blank sum and prc files again. Jay fixed it the issue without having to reboot the pcfs. As part of the same issue, can't send ready/start/finish messages due to error with snp file location (Eloise+Jasper earlier in the week)
  • ohg137: pcfs rebooted to fix prc/snp transfer script again - this time scp wasn't working - manual drudg settings? copy skd, go to previous experiment and copy across check mk5 procedure. Jay to put in notes steps for copying and drudging manually.